• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name O. Glamazdin

    Log entry time 16:06:08 on October17,2006

    Entry number 187413

    keyword=Day shift summary

    8:30 MCC stops beam delivery for injector problems
    8:30 Heidi calls to inform us that the left arm IOC has frozen - we reboot it
    8:30 Sasha notices that some of the ADC's in the right arm have gone down.
    9:10 After rebooting the ROC's, the ADC's come back up.
    10:40 Controlled access for S. Nanda
    10:45 Coda crash
    11:05 Hall A back to Beam permit
    12:26 Beam is back to the Hall A
    12:30 Restart Coda. Reset ROC1-4
    12:50 - 12:55 B5 Production Run 1433 0.4M
    12:58 - 13:20 Run 1434 2M
    13:21 - 13:46 Run 1435 1.9M
    13:47 Half-wave plate is IN
    13:51 - 14:10 Run 1436 1.7M
    14:10 - 16:00 Spot move